OpenStudy (anonymous):

find the geometric mean of 5 and 8 write the exact answer

jimthompson5910 (jim_thompson5910):

geometric mean of x and y = \(\Large \sqrt{x*y}\)

OpenStudy (anonymous):

??? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

multiply the two numbers, then take the square root

OpenStudy (anonymous):

i got 40 but is that the whole answer

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{40}\] now simplify

OpenStudy (anonymous):

im not sure how

jimthompson5910 (jim_thompson5910):

what factors of 40 are perfect squares?

OpenStudy (anonymous):

10?

jimthompson5910 (jim_thompson5910):

4 actually

jimthompson5910 (jim_thompson5910):

since 2 squared = 4

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{40} = \sqrt{4*10}\] \[\Large \sqrt{40} = \sqrt{4}*\sqrt{10}\] \[\Large \sqrt{40} = 2\sqrt{10}\]

OpenStudy (anonymous):

is that the full answer

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you